# Inferring hyperuniformity from local structures via persistent homology

### v1

@@ Abstract
Hyperuniformity refers to the suppression of density fluctuations at large scales. Typical for ordered systems, this property also emerges in several disordered physical and biological systems, where it is particularly relevant to understand mechanisms of pattern formation and to exploit peculiar attributes, e.g., interaction with light and transport phenomena. While hyperuniformity is a global property, it has been shown in [Phys. Rev. Research 6, 023107 (2024)] that global hyperuniform characteristics systematically correlate with topological properties representative of local arrangements. In this work, building on this information, we explore and assess the inverse relationship between hyperuniformity and local structures in point distributions as described by persistent homology. Standard machine learning algorithms trained on persistence diagrams are shown to detect hyperuniformity with high accuracy. Therefore, we demonstrate that the information on patterns’ local structure allows for inferring hyperuniformity. Then, addressing more quantitative aspects, we show that parameters defining hyperuniformity globally, for instance entering the structure factor, can be reconstructed by comparing persistence diagrams of targeted patterns with reference ones. We also explore the generation of patterns entailing given topological properties. The results of this study pave the way for advanced analysis of hyperuniform patterns including local information, and introduce basic concepts for their inverse design.
@@ 1 Introduction
Controlling the design of HU arrangements has also proved useful to tailoring properties in resilient environments as shown, for instance, for some applications in optics [2, 11, 12]. In this context, generating hyperuniform patterns is crucial, and several methods have been proposed [13, 14, 15, 16]. The inverse design of HU structures given a targeted property has been explored less. An interesting aspect of this task consists of the existence of manifolds of HU patterns that are similar in terms of long-range decrease of density fluctuations while exhibiting different arrangements on small scales or, vice versa, they are indistinguishable by the naked eye but globally exhibit different behaviors [2].
Given its generality, hyperuniformity has been widely investigated for arrangements of points and fields of different kinds [2, 17]. It is usually defined in terms of global quantities, such as the number variance sampled over large regions or, equivalently, the limit of the structure factor for vanishing wave numbers. However, although hyperuniformity is usually quantified in terms of global characteristics, it also affects arrangements on relatively small scales. This has been recently shown in reference [18] with the analysis of the statistics of local neighborhood motifs [19, 20] of HU and non-HU point patterns, as well as using an up-and-coming technique which allows the characterization of the “shape” of data sets, the so-called persistent homology [21, 22, 23]. The latter, exploited in this work too, is a topological data analysis method that tracks the appearance (birth) and disappearance (death) of homological features (such as connected components, loops, and voids) upon a given parametrization of patterns (e.g., a threshold for scalar fields) and is further detailed below. Statistics of local neighborhoods and persistent homology unveiled crucial evidence. For instance, different subsets of an ideal HU pattern –resembling realistic conditions realized in a finite-size and non-periodic sampling of experimental systems or numerical simulations– have different structure factors but very similar topological properties. These properties may be then used for their detection and inverse design. However, the latter is yet to be addressed in detail.
@@
This work keeps focusing on the correspondence between hyperuniformity and topological features of HU point arrangements first disclosed in reference [18]. In particular, given the direct connection between global HU character and local arrangements discussed therein, it analyzes the corresponding inverse relations. In other words, we inspect if and how hyperuniformity and specific HU characters can be devised by looking at local arrangements characterized by persistent homology. The latter involves data sets conveniently represented in the so-called persistence diagrams, which cannot be inverted straightforwardly. However, remarkable methods have been proposed that consider representative features, combining the information retained in these persistence diagrams and their exploitation in machine learning (ML) models [24]. As shown below, this enables devising methods to determine whether a given persistent homology corresponds to a HU point arrangement with high accuracy. Moreover, quantifying the (Wasserstein) distance between persistence diagrams with respect to reference ones allows us to estimate the effective structure factor parametrization. The inverse problem of reconstructing patterns from persistent homology is also explored, and we propose a method to achieve it in an approximate manner.
@@
The manuscript is organized as follows. In section 2, we recall the notion of hyperuniformity together with the essential quantities used in this work, as well as details concerning the generation of ideal HU patterns. Section 3 reviews the concept of persistent homology and the related persistence diagrams used to characterize the local arrangements in HU (and non-HU) patterns. The detection of HU or non-HU characters via ML models is reported in section 4. In section 5, we illustrate the estimation of effective structure-factor parametrization by leveraging comparisons between persistence diagrams with reference ones. We finally report on the construction of point arrangements to encode specific information about their geometry (persistence diagrams) in section 6. The main conclusions are discussed and summarized in section 7.
@@
Figure 1: Examples of point patterns with a structure factor as in equation (5). First row: Point arrangements for different structure factor parameters reported above the panels. Second row: Structure factors, illustrated by both the $\mathcal{S}_{0}(|\mathbf{k}|)$ imposed in the generation procedure (solid black lines) and discrete values computed for the patterns in the first rows for $\mathcal{S}(|\mathbf{k}|)$ imposed at appropriate wavenumbers $|\mathbf{k}|=2\pi n$ with $n\in\mathbb{N}_{>0}$ (red dots). We remark that $\widetilde{H}$ in equation (6) corresponds to the value of $\mathcal{S}(|\mathbf{k}|)$ for the $|\mathbf{k}|=2\pi$, i.e. the leftmost red point in the structure factor plots and may deviate significantly from the parameter $H$ for relatively small $\alpha$. See, for instance, the case with $\alpha=0.5$ (first column).
@@ 2 Hyperuniformity
Consider a set $\mathcal{P}$ of points within the Euclidean space $\mathbb{R}^{n}$. For a given spherical observation window with radius $R$, namely a ball $\mathcal{B}_{R}$, we can define the number variance $\sigma(R)\,$=$\,\langle\mathcal{N}(R)^{2}\rangle-\langle\mathcal{N}(R)\rangle^{2}$ with $\mathcal{N}(R)$ the number of points within $\mathcal{B}_{R}$. $\mathcal{P}$ is said to be HU if $\sigma(R)$ increases slower than the volume of $\mathcal{B}_{R}$ [2], i.e.:
$\lim\limits_{R\rightarrow\infty}\frac{\sigma(R)}{R^{n}}=0.$ (1)
@@
In this work, we consider ideal patterns with a prescribed HU character. A point arrangement with a target structure factor $\mathcal{S}_{0}$ can be obtained by minimizing the following objective function [13, 14]
@@
with $\mathcal{S}(\mathbf{k})$ from equation (3) and $Q$ a domain around the origin. Starting from an initial distribution for $\{\mathbf{p}_{j}\}$, the targeted patterns can then be obtained via gradient-based minimization methods with point coordinates as variables. Here, the Broyden–Fletcher–Goldfarb–Shanno (BFGS) minimization algorithm is considered.
@@
In this formula, $\alpha$ controls the power-law scaling of the structure factor for $|\mathbf{k}|\,$$<$$\,K$. We remark that large $\alpha$ allows for approximating the so-called stealthy HU character, i.e. $\mathcal{S}(\mathbf{k})\,$$\sim$ 0 0\,0 for $|\mathbf{k}|\,$$<$$\,K$. $K$ determines the wave numbers for which $\mathcal{S}$ deviates from $1$, i.e., it effectively sets the length scale above which correlations are imposed. For 2D stealthy HU configurations, an order-disorder transition is obtained for $K_{\rm c}\,$=$\,\sqrt{8\pi N}$ [2]. $H$ controls the limit of the structure factor for $|\mathbf{k}|\rightarrow 0$. We recall that this parameter is 00 for an ideal HU pattern of infinite size, as dictated by equation (2). For a finite size pattern, however, only the structure factor for the smallest appropriate $|\mathbf{k}|$ can be computed, namely $\min(|\mathbf{k}|)\,$=$\,k_{\rm min}\,$= $\,2\pi$, which may deviate from 00 even for $H=0$. An effective value for the limit of the structure factor for $|\mathbf{k}|\rightarrow 0$ is thus given by
$\widetilde{H}=\frac{\mathcal{S}(k_{\rm min})}{\mathcal{S}(K)}=\mathcal{S}(2\pi),$ (6)
@@
with the normalization used to generally account for eventual peaks in the structure factor at $K$ [2], while not being relevant for the current parametrization of the structure factor as $\mathcal{S}_{0}(K)\,$= 11\,11 in equation (5). The quantity $\widetilde{H}$ is typically considered in realistic settings, e.g., experiments or numerical simulations, to practically define the HU character. The commonly adopted terminology refers to nearly HU patterns for $H\leq 10^{-2}$ and effective HU patterns for $H\leq 10^{-4}$ [25, 26]. In this article, for discussing geometrical features of HU arrangements, we will consider that a pattern is hyperuniform if $H\leq 10^{-2}$. In [18], it was indeed shown that local arrangements in patterns of the size considered here are not distinguishable for $H<10^{-2}$, motivating our choice. As detailed below, this condition will be practically sampled by means of the accessible $\widetilde{H}$ defined in equation (6).
Figure 2: Illustration of the Čech persistent homology and the persistence diagram for a point pattern of four points. In this case, a parametrization of the simplicial complexes is given by the radius of circles centered in the points, called time $t\in\mathbb{R}$, which realizes the filtration of the point pattern’s homological features. Seven different time values $t_{n}$ with $n=1,\dots,7$ are here presented, resulting in a nested family of seven simplicial complexes. The time at which a homological feature appears is called birth time ($b_{j}$). When it disappears, it is called death time ($d_{j}$). See, for instance, the four connected components (or 0-dimensional holes) at $t_{1}$ born at $t_{0}=0$ and the disappearance of one of them at $t_{2}$ as three connected components can be then identified. On the bottom right, the corresponding persistence diagram is illustrated. It shows points $(b_{j},d_{j})$ obtained with the considered filtration. Different point sets enter the persistence diagram as different homological features can be considered. In 2D, 0- and 1-dimensional holes are the accessible homological features, with the related point sets in the persistence diagram referred to as $h_{0}$ and $h_{1}$, respectively.
@@
To increase the efficiency of the minimization of $O(\{\mathbf{p}_{j}\})$, equation (4), we set as initial configurations slightly (randomly) perturbed triangular lattices (different for every realization). Examples of generated patterns are given in figure 1.
@@ 3 Persistent homology
Persistent homology is a technique used in topological data analysis, leveraging concepts from algebraic topology, to identify and characterize features of data that are persistent across multiple scales [21, 22]. We refer to [23, 27, 28] and reference therein for general and comprehensive treatments of persistent homology while recalling here the main, practical concepts only for point clouds in two spatial dimensions (2D).
@@
This technique typically considers families of simplicial complexes, i.e., collection of points, lines, triangles, tetrahedrons, and their higher-dimensional counterparts, which can be constructed from the considered data set. In particular, families of nested simplicial complexes $X_{t}$ parameterized by a scalar, positive parameter $t$, called time, such that $X_{s}\subseteq X_{t}$ if $s\leq t$ are considered; see figure 2. Topological features, e.g., the number of connected components (or 0-dimensional holes) or 1-dimensional holes, are then evaluated for $X_{t}$ upon varying $t$. Their variation allows for the characterization of the shape of the considered data. For instance, persistent features like a broad region without points within an otherwise uniformly distributed point arrangement could be detected as a peculiar 1-dimensional hole, featured in different $X_{t}$ for a significantly large interval of $t$-values compared to other 1-dimensional holes.
@@
In practice, rules must be specified to construct the simplicial complexes $X_{t}$ while varying $t$ – a procedure called filtration. In this article, extending the preliminary analysis reported in reference [18], two approaches are considered. The first one evaluates the persistent homology exploiting the so-called Čech complex, also equivalent to the alpha complex constructed from the finite cells of a Voronoi tessellation. The second approach exploits the so-called Vietoris-Rips complex [28]. We recall their definitions and further information in A. Both approaches define simplicial complexes from the set of circles with radius $t$ centred on the points in the considered data set; see figure 2 explicitly showing a filtration via Čech complex. For each $k$-dimensional hole ($k=0,1$ in 2D) in the simplicial complexes obtained by varying $t$, the birth time $b^{k}$ (appearance) and the death time $d^{k}$ (disappearance) are determined.
@@
A typical representation of these quantities is the so-called persistence diagram, which considers points $\mathbf{p}_{j}^{k}\,$=$\,(b^{k}_{j},d^{k}_{j})$ in the corresponding Cartesian representation with birth and death times as $x$ and $y$ coordinates, also illustrated in figure 2. Note that $b^{0}_{j}\,$= 0 0\,0, as connected components (0-dimensional holes) have a birth time of $t\,$= 0 0\,0 by definition. The points $\mathbf{p}_{j}^{k}$ characterizing the persistent homology and entering persistence diagrams are hereafter called ${h}_{k}$ points and their collection $h_{k}$-diagram for brevity. Further discussions and illustrations can be found in reference [24].

### v2

# Inferring traits of hyperuniformity from local structures via persistent homology

@@ Abstract
Hyperuniformity refers to the suppression of density fluctuations at large scales. Typical for ordered systems, this property also emerges in several disordered physical and biological systems, where it is particularly relevant to understand mechanisms of pattern formation and to exploit peculiar attributes, e.g., interaction with light and transport phenomena. While hyperuniformity is a global property, ideally defined for infinitely extended systems, several disordered correlated systems have finite size. It has been shown in [Phys. Rev. Research 6, 023107 (2024)] that global hyperuniform characteristics systematically correlate with distributions of topological properties representative of local arrangements. In this work, building on this information, we explore and assess the inverse relationship between hyperuniformity and local structures in point patterns as described by persistent homology. Standard machine learning algorithms trained on persistence diagrams are shown to detect hyperuniformity of periodic point patterns with high accuracy. Therefore, we demonstrate that the information on patterns’ local structures allows for characterizing whether finite size arrangements are analogous to those realized in hyperuniform patterns. Then, addressing more quantitative aspects, we show that parameters defining hyperuniformity globally can be reconstructed by comparing persistence diagrams of targeted patterns with reference ones. We also explore the generation of patterns entailing given topological properties. The results of this study pave the way for advanced analysis of hyperuniform patterns including local information, and introduce basic concepts for their inverse design.
@@ 1 Introduction
Controlling the design of HU arrangements has also proved useful to tailoring properties in resilient environments as shown, for instance, for some applications in optics [2, 11, 12]. In this context, generating hyperuniform patterns is crucial, and several methods have been proposed [13, 14, 15, 16, 17]. The inverse design of HU structures given a targeted property has been explored less. Recently, an approach to inverse design HU, typical non-HU and anti-HU patterns by tuning their pair statistics has been proposed [18]. An interesting aspect of this task consists of the existence of manifolds of HU patterns that are similar in terms of long-range decrease of density fluctuations while exhibiting different arrangements on small scales or, vice versa, they are indistinguishable by the naked eye but globally exhibit different behaviors [2].
Given its generality, hyperuniformity has been widely investigated for arrangements of points and fields of different kinds [2, 19]. It is usually defined in terms of global quantities, such as the number variance sampled over large regions or, equivalently, the limit of the structure factor for vanishing wave numbers. However, although hyperuniformity is usually quantified in terms of global characteristics, it also affects arrangements on relatively small scales. In addition, physical systems that may exhibit hyperuniformity are very often realized over finite-size domains. This is, for instance, the case of metamaterials for mechanical applications leveraging spinodal patterns [20]. It may also result from the self-assembly of HU patterns [10] when confined to prescribed finite-size regions (e.g., via patterning), and could emerge in HU active systems [21] which are often confined to one phase in two- or multi-phase systems [22]. Descriptions beyond classical analysis of long-wavelength behavior are thus in general needed to fully characterize these systems.
Correlations between hyperuniformity and distributions of local structures have been recently shown in reference [23] with the analysis of the statistics of local (graph) neighborhood motifs [24, 25] of HU and non-HU point patterns, as well as using an up-and-coming technique which allows the characterization of the “shape” of data sets, the so-called persistent homology [26, 27, 28]. The latter, exploited in this work too, is a topological data analysis method that tracks the appearance (birth) and disappearance (death) of homological features (such as connected components, loops, and voids) upon a given parametrization of patterns (e.g., a threshold for scalar fields) and is further detailed below. Importantly, persistent homology analyzes patterns via a parametrization that echoes characterizations and constructions already discussed in the context of hyperuniformity for two-phase media, such as the decoration of points with circles [29, 2] and the concept of spreadability of a solute in a surrounding matrix [30].
Statistics of local neighborhoods and persistent homology applied to HU patterns unveiled crucial evidence. For instance, different subsets of an ideal HU pattern –resembling realistic conditions realized in a finite-size and non-periodic sampling of experimental systems or numerical simulations– have different structure factors but similar distributions of topological properties. These may be then used for their detection and inverse design. However, except first attempts [18], the latter is yet to be addressed in detail.
@@
This work keeps focusing on the correspondence between hyperuniformity and topological features of HU point patterns first disclosed in reference [23]. In particular, given the direct connection between global HU character and local arrangements discussed therein, it analyzes the corresponding inverse relations. In other words, we inspect if and how hyperuniformity and specific HU characters can be devised by looking at local arrangements characterized by persistent homology. More in general, we aim to determine if geometrical arrangements in a given pattern are similar to those realized in ideal HU patterns. Persistent homology involves data sets conveniently represented in the so-called persistence diagrams, which cannot be inverted straightforwardly. However, remarkable methods have been proposed that consider representative features, combining the information retained in these persistence diagrams and their exploitation in machine learning (ML) models [31]. As shown below, this enables devising methods to determine whether a given persistent homology corresponds to a HU point pattern with high accuracy. Moreover, quantifying the (Wasserstein) distance between persistence diagrams with respect to reference ones allows us to estimate the effective structure factor parametrization. The inverse problem of reconstructing patterns from persistent homology is also explored, and we propose a method to achieve it in an approximate manner.
@@
The manuscript is organized as follows. In section 2, we recall the notion of hyperuniformity together with the essential quantities used in this work, as well as details concerning the generation of ideal HU patterns. Section 3 reviews the concept of persistent homology and the related persistence diagrams used to characterize the local arrangements in HU (and non-HU) patterns. The detection of HU or non-HU characters via ML models is reported in section 4. In section 5, we illustrate the estimation of effective structure-factor parametrization by leveraging comparisons between persistence diagrams with reference ones. Section 6 additionally shows that the information contained in persistence diagrams can be exploited to detect boundaries between different realizations of nominally HU patterns. We finally report on the construction of point patterns to encode specific information about their geometry (persistence diagrams) in section 7. The main conclusions are discussed and summarized in section 8. The data and scripts used in this study are available at [32].
@@
Figure 1: Examples of point patterns with a structure factor as in equation (5). First row: Point patterns for different structure factor parameters reported above the panels. Second row: Structure factors, illustrated by both the $\mathcal{S}_{0}(|\mathbf{k}|)$ imposed in the generation procedure (solid black lines) and discrete values computed for the patterns in the first rows for $\mathcal{S}(|\mathbf{k}|)$ imposed at appropriate wavenumbers $|\mathbf{k}|=2\pi n$ with $n\in\mathbb{N}_{>0}$ (red dots). We remark that $\widetilde{H}$ in equation (6) corresponds to the value of $\mathcal{S}(|\mathbf{k}|)$ for the $|\mathbf{k}|=2\pi$, i.e. the leftmost red point in the structure factor plots and may deviate significantly from the parameter $H$ for relatively small $\alpha$. See, for instance, the case with $\alpha=0.5$ (first column).
@@ 2 Hyperuniformity
Consider a set $\mathcal{P}$ of points within the Euclidean space $\mathbb{R}^{n}$. For a given spherical observation window with radius $R$, namely a ball $\mathcal{B}_{R}$, we can define the number variance $\sigma^{2}(R)\,$=$\,\langle\mathcal{N}(R)^{2}\rangle-\langle\mathcal{N}(R)\rangle^{2}$ with $\mathcal{N}(R)$ the number of points within $\mathcal{B}_{R}$. $\mathcal{P}$ is said to be HU if $\sigma^{2}(R)$ increases slower than the volume of $\mathcal{B}_{R}$ [2], i.e.:
$\lim\limits_{R\rightarrow\infty}\frac{\sigma^{2}(R)}{R^{n}}=0.$ (1)
@@
In this work, we consider ideal patterns with a prescribed HU character. A point pattern with a target structure factor $\mathcal{S}_{0}$ can be obtained by minimizing the following objective function [13, 14]
@@
with $\mathcal{S}(\mathbf{k})$ from equation (3) and $Q$ a domain around the origin. Starting from an initial distribution for $\{\mathbf{p}_{j}\}$, the targeted patterns can then be obtained via gradient-based minimization methods with point coordinates as variables. Here, the Broyden–Fletcher–Goldfarb–Shanno (BFGS) minimization algorithm is considered. Efficient approaches to numerically solve this problem can be found in Refs. [16, 17].
@@
In this formula, $\alpha$ controls the power-law scaling of the structure factor for $|\mathbf{k}|\,$$<$$\,K$. Large $\alpha$ allows for approximating the so-called stealthy HU character, i.e. $\mathcal{S}(\mathbf{k})\,$$\sim$ 0 0\,0 for $|\mathbf{k}|\,$$<$$\,K$. $K$ determines the wave numbers for which $\mathcal{S}$ deviates from $1$, i.e., it effectively sets the length scale above which correlations are imposed. For 2D stealthy HU configurations, an order-disorder transition is obtained for $K_{\rm c}\,$=$\,\sqrt{8\pi N}$ [2]. $H$ controls the limit of the structure factor for $|\mathbf{k}|\rightarrow 0$.
The structure factor in Eq. (5) is chosen as it features a simple, straightforward parameterization while enforcing global HU characters and well-defined and comparable small-scale randomness. The latter follows from choosing $\mathcal{S}_{0}(\mathbf{k})=1$ for $|\mathbf{k}|\,$$>$$\,K$. Other forms for $\mathcal{S}_{0}(\mathbf{k})$ other than (5) can be considered. In particular, several HU systems exhibit oscillations for $|\mathbf{k}|\,$$>$$\,K$ [2]. Modeling these cases, however, requires introducing additional choices concerning the functional form and parameters while not introducing any qualitative novel aspect. Another approach may consist of not enforcing any value of $\mathcal{S}_{0}(\mathbf{k})$ for $|\mathbf{k}|\,$$>$$\,K$.
We recall that $H=0$ for an ideal HU pattern of infinite size, as dictated by equation (2). For a finite size pattern, however, only the structure factor for the smallest appropriate $|\mathbf{k}|$ can be computed, namely $\min(|\mathbf{k}|)\,$=$\,k_{\rm min}$, which may deviate from 00 even for $H=0$. An effective value for the limit of the structure factor for $|\mathbf{k}|\rightarrow 0$ is thus given by
$\widetilde{H}=\frac{\mathcal{S}(k_{\rm min})}{\mathcal{S}(K)}.$ (6)
@@
with the normalization used to generally account for eventual peaks in the structure factor at $K$ [2]. For the parametrization used in this work, $\widetilde{H}\,$=$\,\mathcal{S}(2\pi)$ as $k_{\rm min}\,$= $\,2\pi$ being the domain a unit square and $\mathcal{S}_{0}(K)\,$= 11\,11. The quantity $\widetilde{H}$ is typically considered in realistic settings, e.g., experiments or numerical simulations, to practically define the HU character. The commonly adopted terminology refers to nearly HU patterns for $H\leq 10^{-2}$ and effective HU patterns for $H\leq 10^{-4}$ [33, 34]. In this article, for discussing geometrical features of HU arrangements, we will consider that a pattern is hyperuniform if $H\leq 10^{-2}$. In [23], it was indeed shown that local arrangements in patterns of the size considered here are not or barely distinguishable for $H<10^{-2}$, motivating our choice. As detailed below, this condition will be practically sampled by means of the accessible $\widetilde{H}$ defined in equation (6).
Figure 2: Illustration of the Čech persistent homology and the persistence diagram for a point pattern of four points.
@@
Any initial arrangement can be, in principle, considered for the minimization of $O(\{\mathbf{p}_{j}\})$, equation (4). An efficient approach consists of setting as initial configurations slightly (randomly) perturbed triangular lattices (different for every realization). This choice allows for faster convergence to the desired pattern compared to initial fully random arrangements. We note that the perturbation is needed to escape the unstable equilibrium represented by periodic lattices. Examples of generated patterns are given in figure 1.
@@ 3 Persistent homology
Persistent homology is a technique used in topological data analysis, leveraging concepts from algebraic topology, to identify and characterize features of data that are persistent across multiple scales [26, 27]. We refer to [28, 35, 36] and reference therein for general and comprehensive treatments of persistent homology while recalling here the main, practical concepts only for point patterns in two spatial dimensions (2D).
@@
This technique typically considers families of simplicial complexes, i.e., collection of points, lines, triangles, tetrahedrons, and their higher-dimensional counterparts, which can be constructed from the considered data set. In particular, families of nested simplicial complexes $X_{t}$ parameterized by a scalar, positive parameter $t$, called time, such that $X_{s}\subseteq X_{t}$ if $s\leq t$ are considered; see figure 2. Topological features, e.g., the number of connected components (or 0-dimensional holes) or 1-dimensional holes, are then evaluated for $X_{t}$ upon varying $t$. Their variation allows for the characterization of the shape of the considered data. For instance, persistent features like a broad region without points within an otherwise uniformly distributed point pattern could be detected as a peculiar 1-dimensional hole, featured in different $X_{t}$ for a significantly large interval of $t$-values compared to other 1-dimensional holes.
@@
In practice, rules must be specified to construct the simplicial complexes $X_{t}$ while varying $t$, a procedure called filtration. In this article, extending the preliminary analysis reported in reference [23], two approaches are considered. The first one evaluates the persistent homology exploiting the so-called Čech complex, also equivalent to the alpha complex constructed from the finite cells of a Voronoi tessellation. The second approach exploits the so-called Vietoris-Rips complex [36]. We recall their definitions and further information in A. Both approaches define simplicial complexes from the set of circles with radius $t$ centered on the points in the considered data set; see figure 2 explicitly showing a filtration via Čech complex for a pattern comprising four points. In particular, this figure shows seven different time values $t_{n}$ with $n=1,\dots,7$, realizing a nested family of seven simplicial complexes. The time at which a homological feature appears is called birth time ($b$). When it disappears, it is called death time ($d$). See, for instance, the four connected components (or 0-dimensional holes) at $t_{1}$ born at $t_{0}=0$ and the disappearance of one of them at $t_{2}$ as three connected components can be then identified.
@@
For each $k$-dimensional hole ($k=0,1$ in 2D) in the simplicial complexes obtained by varying $t$, the birth time $b^{k}$ (appearance) and the death time $d^{k}$ (disappearance) are thus determined. A typical representation of these quantities is the so-called persistence diagram, which considers points $\mathbf{p}_{j}^{k}\,$=$\,(b^{k}_{j},d^{k}_{j})$ in the corresponding Cartesian representation with birth and death times as $x$ and $y$ coordinates, also illustrated in figure 2 (bottom right). Note that $b^{0}_{j}\,$= 0 0\,0, as connected components (0-dimensional holes) have a birth time of $t\,$= 0 0\,0 by definition. The points $\mathbf{p}_{j}^{k}$ characterizing the persistent homology and entering persistence diagrams are hereafter called ${h}_{k}$ points and their collection $h_{k}$-diagram for brevity. Further discussions and illustrations can be found in reference [31].
Persistent homology and persistence diagrams relate to a few concepts well discussed in the context of hyperuniformity. Decorating the points with spheres, similar to filtration, is an approach often used to generate two-phase media (mimicking particles or voids) [29, 2] exploited to study HU phases. Also, the filtration itself can be related to the concept of spreadability in a two-phase medium, namely the diffusion from one phase into a second phase, which has been found to correlate with HU characters [30]. We remark that persistence diagrams contain information concerning the geometry of the analyzed pattern. Information on neighborhoods of all the points is retained well beyond two-body information. Moreover, they have a close relationship to the bounded hole property, which all stealthy HU systems and some non-stealthy HU have [37, 38, 39].